# A Tutorial on Learning-Based Radio Map Construction: Data, Paradigms, and Physics-Awareness

### 18 Mar 2026 v1

@@ IV-D Dataset Analysis and Benchmarks
RadioMapSeer [69] is the foundational benchmark for 2D path loss prediction. It comprises 700 city maps with propagation simulated using both DPM and IRT engines at 10 m resolution. As the most widely used benchmark, most methods in Sections V and VI report results on this dataset. Its main limitation is the absence of vertical information and multipath parameters.
@@
IndoorRadioMap [82] was released for the First Indoor Pathloss RM Prediction Challenge. It includes 25 indoor environments with material properties encoded into RGB tensors at 0.25 m resolution. The simulation considers up to 8 reflections and 10 transmissions with directional antenna patterns.
@@ V Learning-Based Methods for Source-Aware Radio Mapping
Figure 3: The progression illustrates a paradigm shift towards increasing spatial and physical modeling capabilities. Early CNNs rely on local receptive fields but struggle with long-range diffraction. ViTs overcome this via global self-attention to capture macro-shadowing dependencies. GNNs embed physics-informed priors by modeling non-Euclidean multipath topologies. Recently, generative modelsreframe the task as probabilistic inference, enabling the reconstruction of high-fidelity radio maps from extreme sparsity while overcoming physical hallucinations.
This section reviews five neural architecture families for source-aware RM construction. Table VI provides a cross-architecture comparison to guide practitioners. No single architecture dominates across all dimensions [146, 117]. Each family occupies a distinct niche in the trade-off space among latency, data efficiency, spatial modeling, and output modality.
@@ V-A CNN-Based Methods
PPNet [55] SegNet (classification) $H\!\times\!W\!\times\!3$ (Layout, Gauss, Dist) $H\!\times\!W\!\times\!1$ 30 GHz RT outdoor MAE: 1.08 dB N/A
@@ V-A1 Encoder-Decoder Paradigm and Resolution Bottlenecks
PPNet [55] addresses this bottleneck through reformulation. Instead of continuous regression, it quantizes the output into discrete bins and uses cross-entropy loss. Continuous regression over a 100+ dB dynamic range is sensitive to outliers and gradient variance. Discrete classification establishes clear decision boundaries that stabilize training. Although quantization introduces up to 0.5 dB error per bin, the improved robustness outweighs this cost. PPNet also memorizes max-pooling indices during encoding for reuse during decoding, preserving sharp edges without additional parameters.
@@ V-A2 Receptive Field Expansion
UNetDCN [56] integrates deformable convolutions that learn positional offsets for each sampling point. The receptive field expands non-uniformly to selectively attend to distant diffraction edges. A secondary capability is that the trained network can serve as a differentiable surrogate for antenna parameter optimization via backpropagation [107]. However, gradient-based inversion through deep networks is susceptible to gradient explosion and local optima.
PMNet [66] uses atrous convolutions and spatial pyramid pooling (SPP). Atrous convolutions increase the effective receptive field by inserting gaps between kernel elements [65]. The SPP module aggregates multi-scale features at the bottleneck. This captures building occlusion patterns at multiple spatial extents simultaneously. A transfer learning framework further enables generalization across city layouts.
@@ V-B ViT-Based Methods
Indoor propagation introduces further complexity. TransPathNet [73] uses a two-stage coarse-to-fine residual learning framework. A TransNeXt encoder predicts a coarse global map, which is then refined with physical priors in the second stage. DINOv2-ViT [75] leverages large-scale foundation model pre-training for cross-condition generalization. The massive self-supervised pre-training of DINOv2 provides rich embeddings learned from diverse visual structures. This appears to be the dominant contributor to generalization, as evidenced by the gap between fine-tuned DINOv2 and randomly initialized architectures.
@@
Tutorial takeaway: ViTs offer the greatest advantage when inputs are spatially sparse and long-range blockage correlations exist [74, 34]. When inputs are dense and the environment is locally governed, the performance gap narrows. The foundation model paradigm represents a promising direction for cross-environment generalization [75, 138].
@@ V-D GAN-Based Methods
$\min_{G}\max_{D}\;$ $\mathbb{E}_{x\sim P_{data}}[\log D(Q\mid u)]$
$+\mathbb{E}_{z\sim P_{z}}[\log(1-D(G(P\mid u)))].$ (29)
@@ V-D2 Unpaired and Unconditional Generation
CycleGAN [101] eliminates pixel-to-pixel correspondence using cycle-consistency and identity losses:
@@ VI-A CNN-Based Methods
CAED [115] 3D conv. autoencoder Extreme 3D 3D voxel Decoder dilation + structural pruning 3D Frobenius
@@ VI-B ViT-Based Methods
The Transformer structure restoration (TSR) framework [21] employs alternating self-attention and axial attention modules. This reduces complexity from $O(n^{2})$ to $O(2n^{3/2})$. Standard $L_{2}$ loss averages multiple plausible patterns, yielding blurred reconstructions. TSR addresses this via a high-receptive field (HRF) perceptual loss using a pre-trained segmentation network:
@@
where $U$ is the rotation matrix aligning with $v(r)$. If the entire environment and transmitter are translated or rotated, the RM transforms accordingly. By constructing features in a transmitter-relative frame, STORM guarantees this equivariance structurally [46]. This eliminates the need for data augmentation to learn basic spatial symmetries. STORM extends to active sensing via causal self-attention for joint estimation and next-best-point selection.
@@ VI-D Diffusion-Based Methods
Deterministic models output a single point estimate minimizing mean squared error [69, 66]. However, this is insufficient when downstream applications require nonlinear RM functionals [42, 83]. Consider an NLoS location where RSS fluctuates between 85-85 and 95-95 dBm. A deterministic model predicts 90-90 dBm. Channel capacity $C=B\log_{2}(1+\text{SNR})$ is concave in linear-scale power. By Jensen’s inequality, the point estimate systematically overestimates expected capacity [42]. This bias can reach several bits/s/Hz in low-SNR regimes. Diffusion models resolve this by generating $J$ samples from $p(\gamma|\mathcal{M})$ and computing unbiased Monte Carlo estimates: $\hat{\phi}_{B}\approx\frac{1}{J}\sum_{j=1}^{J}F(\gamma_{j})$ [83].
@@ VII Optics-Inspired Electromagnetic Mapping Methods
The methods in Sections V and VI treat RM construction as image-level regression. This section reviews an alternative paradigm inspired by neural rendering [143, 62, 147]. NeRF and 3DGS construct continuous, physics-grounded representations by explicitly modeling the volumetric propagation process. Table XV identifies when optics-inspired methods offer genuine advantages over traditional neural networks.
@@ VIII-B1 Material and Antenna Properties
$\text{FSPL}+G_{2D}=20\log_{10}\!\left(\frac{4\pi Df}{c_{o}}\right)+G(\Phi_{2D},\Theta_{2D}).$ (52)
@@ VIII-B4 Topological Filtering for Dynamic Environments
Dynamic environments with moving sources introduce multipath artifacts that corrupt the input feature space [39, 150]. Topological data analysis (TDA) distinguishes genuine sources from transient reflections via the concept of persistence. A genuine radiation source produces a broad peak that persists across a wide threshold range. A multipath artifact produces a narrow peak with low persistence. Formally, the 0-th persistent homology group assigns each local maximum a birth and death level. Peaks with persistence below a physical threshold are filtered as artifacts [118]. The computational complexity is $O(n^{2})$ to $O(n^{3})$, where $n$ is the grid point count. For a $256\times 256$ grid, this can require several seconds on standard hardware.
@@ VIII-C4 Helmholtz Equation Embedding
The first component $\mathcal{L}_{PDE}$ penalizes violations of the Helmholtz equation across the entire spatial domain. It evaluates the PDE residual at every grid point by computing the discrete Laplacian of the predicted field $\hat{u}$ and checking consistency with the local wavenumber and source term [54, 59]. This residual-based penalty constrains the predicted field to conform to wave propagation dynamics even in regions lacking measurement samples. The second component $\mathcal{L}_{BC}$ enforces Dirichlet boundary conditions at the interfaces between free space and physical obstacles [54]. At building walls and structural boundaries, the electromagnetic field must satisfy material-dependent boundary constraints. By penalizing deviations from these boundary values, $\mathcal{L}_{BC}$ ensures that the predicted field correctly reflects the attenuation and reflection behavior imposed by the environmental geometry [1]. The third component $\mathcal{L}_{Source}$ constrains the predicted source term at the known transmitter location [54]. It anchors the radiation origin by penalizing discrepancies between the predicted field at the source node and the theoretical emission pattern. This prevents the network from generating physically plausible wave fields that originate from incorrect spatial locations. The three components address complementary physical aspects: $\mathcal{L}_{PDE}$ governs wave propagation in free space, $\mathcal{L}_{BC}$ governs wave-obstacle interaction at boundaries, and $\mathcal{L}_{Source}$ governs the radiation origin. Together, they force the latent representations extracted under ultra-sparse observations to strictly conform to wave propagation dynamics [54]. This discretization introduces a truncation error of $\mathcal{O}(\Delta x^{2})$ [33]. At 28 GHz with $\lambda\approx 10.7$ mm and grid resolution $\Delta x=1$ m, the fourth derivative $\partial^{4}u/\partial x^{4}$ scales as $(2\pi/\lambda)^{4}\approx 1.2\times 10^{9}$ m-4. The resulting truncation error can exceed the actual path loss variation per cell. In practice, higher-order finite difference schemes or multi-scale refinements are needed [111]. Once properly discretized, the combined regularization $\mathcal{L}_{PLN}$ yields a 37.2% NMSE improvement at 1–10% sampling rates over diffusion-only baselines [54].
@@ VIII-D3 Physics-Informed Generative Architectures
VIII-D4 Physical Hallucinations: Definition and Detection
VIII-E Summary and Practitioner Guidance
@@ IX Open Issues and Future Directions
Learning-based RM construction has achieved significant milestones [146, 117]. However, the transition toward ubiquitous 6G environmental awareness remains constrained by several open challenges. These span two axes: modeling of uncharted propagation scenarios and algorithmic frontiers in inverse problems and physics-data co-design. This section identifies the most pressing issues, drawing upon the architectural insights, physical embedding strategies, and dataset limitations discussed throughout this tutorial.
@@ X Concluding Remarks
This tutorial has systematically surveyed the data foundations, neural architectures, and physics-informed strategies for learning-based RM construction. A central thesis emerges: the progression from purely data-driven models toward physics-data dual-driven frameworks is a necessity dictated by the information-theoretic structure of the problem. When observations are sparse and propagation is complex, data alone cannot uniquely determine the RM. The model must draw upon physical priors to resolve ambiguity. The three levels of physics integration surveyed in Section VIII provide progressively deeper embedding of electromagnetic knowledge. Input-level feature engineering offers the lowest entry barrier. Loss-level PDE regularization constrains solutions to feasible electromagnetic manifolds. Architecture-level structural isomorphism achieves the deepest alignment with wave propagation dynamics. Each level trades computational cost for physical fidelity. Several converging trends point toward the future of the field. Differentiable ray tracing enables joint simulation-optimization. Foundation model pre-training supports cross-environment generalization. Amortized inference can bring real-time rendering to optics-inspired methods. Formal hallucination detection provides deployment assurance. Together, these trends suggest a future where RM construction transitions from an offline, per-scenario engineering task to an always-on, physically rigorous component of the wireless network infrastructure. Achieving this vision requires sustained collaboration between the wireless communications, computational electromagnetics, and machine learning communities. The most powerful RM models will be those that respect the physics they seek to represent.

### 28 Mar 2026 v4

@@ IV-D Dataset Analysis and Benchmarks
RadioMapSeer [73] is the foundational benchmark for 2D path loss prediction. It comprises 701 city maps with 80 transmitter positions per map, yielding a total of 56,080 radio maps simulated using both DPM and IRT engines at 5.9 GHz with 1 m resolution. As the most widely used benchmark, most methods in Sections V and VI report results on this dataset. Its main limitation is the absence of vertical information and multipath parameters.
@@
IndoorRadioMap [10] was released for the First Indoor Pathloss RM Prediction Challenge [9]. It includes 25 indoor environments with material properties encoded into RGB tensors at 0.25 m resolution. The simulation considers up to 8 reflections and 10 transmissions with directional antenna patterns.
@@ V Learning-Based Methods for Source-Aware Radio Mapping
Figure 4: The progression illustrates a paradigm shift towards increasing spatial and physical modeling capabilities. Early CNNs rely on local receptive fields but struggle with long-range diffraction. ViTs overcome this via global self-attention to capture macro-shadowing dependencies. GNNs embed physics-informed priors by modeling non-Euclidean multipath topologies. Recently, generative models reframe the task as probabilistic inference, enabling the reconstruction of high-fidelity radio maps from extreme sparsity. When combined with physics-informed constraints, these models can further mitigate physical hallucinations.
This section reviews five neural architecture families for source-aware RM construction. Table VI provides a cross-architecture comparison to guide practitioners. No single architecture dominates across all dimensions [152, 120]. Each family occupies a distinct niche in the trade-off space among latency, data efficiency, spatial modeling, and output modality. Table VIII further catalogs reusable architectural modules that can be integrated into these families without fundamental redesign. Fig. 4 illustrates the architectural progression from local convolutional receptive fields to global generative inference.
@@ V-A CNN-Based Methods
PPNet [100] SegNet (regression) $256\!\times\!256\!\times\!2$ (Height map, Tx) $256\!\times\!256\!\times\!1$ RadioMapSeer (5.9 GHz) RMSE: 0.0507 (norm.) N/A
@@ V-A1 Encoder-Decoder Paradigm and Resolution Bottlenecks
PPNet [100] addresses this bottleneck through architectural design inherited from SegNet. The pooling layers store the indices of maximum activations during encoding and reuse them during decoding. This preserves fine-grained spatial information related to building geometry without additional parameters. A two-phase training strategy first learns binary building layouts, then fine-tunes with height-encoded inputs at a reduced learning rate. This curriculum stabilizes convergence and improves accuracy near building boundaries.
@@ V-A2 Receptive Field Expansion
UNetDCN [59] integrates deformable convolutions that learn positional offsets for each sampling point. The receptive field expands non-uniformly to selectively attend to distant diffraction edges. A secondary capability is that the trained network can serve as a differentiable surrogate for antenna parameter optimization via backpropagation. However, gradient-based inversion through deep networks is susceptible to gradient explosion and local optima.
PMNet [70, 69] uses atrous convolutions and spatial pyramid pooling (SPP). Atrous convolutions increase the effective receptive field by inserting gaps between kernel elements [69]. The SPP module aggregates multi-scale features at the bottleneck. This captures building occlusion patterns at multiple spatial extents simultaneously. A transfer learning framework further enables generalization across city layouts.
@@ V-B ViT-Based Methods
Indoor propagation introduces further complexity. TransPathNet [77] uses a two-stage coarse-to-fine residual learning framework. A TransNeXt encoder predicts a coarse global map, which is then refined with physical priors in the second stage. DINOv2-ViT [90] leverages large-scale foundation model pre-training for cross-condition generalization. It employs a DINOv2 vision transformer as the encoder and a UPerNet convolutional decoder for reconstruction. The massive self-supervised pre-training of DINOv2 provides rich embeddings learned from diverse visual structures. Cross-condition evaluation across unseen buildings, frequencies, and antenna configurations demonstrates the robustness of this transfer learning approach.
@@
Tutorial takeaway: ViTs offer the greatest advantage when inputs are spatially sparse and long-range blockage correlations exist [78, 37]. When inputs are dense and the environment is locally governed, the performance gap narrows. The foundation model paradigm represents a promising direction for cross-environment generalization [79].
@@ V-D GAN-Based Methods
$\min_{G}\max_{D}\;$ $\mathbb{E}_{x\sim p_{\mathrm{data}}}[\log D(x\mid u)]$
$+\mathbb{E}_{z\sim p_{z}}[\log(1-D(G(z\mid u)))].$ (29)
@@ V-D2 Unpaired and Unconditional Generation
CycleGAN [169] eliminates pixel-to-pixel correspondence using cycle-consistency and identity losses. Roger et al. [104] adapt this framework to V2X radio map reconstruction:
@@ VI-A CNN-Based Methods
CAED [52] 3D conv. autoencoder 1–95% 3D 3D voxel Decoder dilation + structural pruning 3D Frobenius
@@ VI-B ViT-Based Methods
The Transformer structure restoration (TSR) framework [23] employs alternating self-attention and axial attention modules. This reduces complexity from $O(n^{2})$ to $O(n^{3/2})$. Standard $L_{2}$ loss averages multiple plausible patterns, yielding blurred reconstructions. TSR addresses this via a high-receptive field (HRF) perceptual loss using a pre-trained segmentation network:
@@
where $U$ is the rotation matrix aligning with $v(r)$. If the entire environment and transmitter are translated or rotated, the RM transforms accordingly. By constructing features in a query-point-relative frame, STORM guarantees this equivariance structurally [121]. This eliminates the need for data augmentation to learn basic spatial symmetries. STORM extends to active sensing via causal self-attention for joint estimation and next-best-point selection.
@@
Figure 5: Illustration of the forward noising process and the reverse denoising process in diffusion-based radio map reconstruction.
@@ VI-D Diffusion-Based Methods
Deterministic models output a single point estimate minimizing mean squared error [73, 70]. However, this is insufficient when downstream applications require nonlinear RM functionals [45, 86]. Consider an NLoS location where RSS fluctuates between 85-85 and 95-95 dBm. A deterministic model predicts 90-90 dBm. Channel capacity $C=B\log_{2}(1+\text{SNR})$ is concave in linear-scale power. By Jensen’s inequality, the point estimate systematically overestimates expected capacity [45]. This bias can reach several bits/s/Hz in low-SNR regimes. Diffusion models resolve this by generating $J$ samples from $p(\gamma|\mathcal{M})$ and computing unbiased Monte Carlo estimates: $\hat{\phi}_{B}\approx\frac{1}{J}\sum_{j=1}^{J}F(\gamma_{j})$ [86]. Fig. 5 visualizes the forward noising and reverse denoising processes in diffusion-based radio map reconstruction.
@@ VII Optics-Inspired Electromagnetic Mapping Methods
The methods in Sections V and VI treat RM construction as image-level regression. This section reviews an alternative paradigm inspired by neural rendering [161, 66, 153, 141]. NeRF and 3DGS construct continuous, physics-grounded representations by explicitly modeling the volumetric propagation process. Table XV identifies when optics-inspired methods offer genuine advantages over traditional neural networks.
@@ VIII-B1 Material and Antenna Properties
$\text{FSPL}+G_{2D}=20\log_{10}\!\left(\frac{4\pi Df}{c}\right)+G(\Phi_{2D},\Theta_{2D}).$ (52)
@@ VIII-B4 Topological Filtering for Dynamic Environments
Dynamic environments with moving sources introduce multipath artifacts that corrupt the input feature space [42, 157]. Topological data analysis (TDA) distinguishes genuine sources from transient reflections via the concept of persistence. A genuine radiation source produces a broad peak that persists across a wide threshold range. A multipath artifact produces a narrow peak with low persistence. Formally, the 0-th persistent homology group assigns each local maximum a birth and death level. Peaks with persistence below a physical threshold are filtered as artifacts. The computational complexity is $O(n^{2})$ to $O(n^{3})$, where $n$ is the grid point count. For a $256\times 256$ grid, this can require several seconds on standard hardware.
@@ VIII-C4 Helmholtz Equation Embedding
The first component $\mathcal{L}_{PDE}$ penalizes violations of the Helmholtz equation across the entire spatial domain. It evaluates the PDE residual at every grid point by computing the discrete Laplacian of the predicted field $\hat{u}$ and checking consistency with the local wavenumber and source term [56, 63]. This residual-based penalty constrains the predicted field to conform to wave propagation dynamics even in regions lacking measurement samples. The second component $\mathcal{L}_{BC}$ enforces Dirichlet boundary conditions at the interfaces between free space and physical obstacles [56]. At building walls and structural boundaries, the electromagnetic field must satisfy material-dependent boundary constraints. By penalizing deviations from these boundary values, $\mathcal{L}_{BC}$ ensures that the predicted field correctly reflects the attenuation and reflection behavior imposed by the environmental geometry [1]. The third component $\mathcal{L}_{Source}$ constrains the predicted source term at the known transmitter location [56]. It anchors the radiation origin by penalizing discrepancies between the predicted field at the source node and the theoretical emission pattern. This prevents the network from generating physically plausible wave fields that originate from incorrect spatial locations. The three components address complementary physical aspects: $\mathcal{L}_{PDE}$ governs wave propagation in free space, $\mathcal{L}_{BC}$ governs wave-obstacle interaction at boundaries, and $\mathcal{L}_{Source}$ governs the radiation origin. Together, they force the latent representations extracted under ultra-sparse observations to strictly conform to wave propagation dynamics [56]. This discretization introduces a truncation error of $\mathcal{O}(\Delta x^{2})$ [36]. At 28 GHz with $\lambda\approx 10.7$ mm and grid resolution $\Delta x=1$ m, the fourth derivative $\partial^{4}u/\partial x^{4}$ scales as $(2\pi/\lambda)^{4}\approx 1.19\times 10^{11}\,\mathrm{m}^{-4}$. The resulting truncation error can exceed the actual path loss variation per cell. In practice, higher-order finite difference schemes or multi-scale refinements are needed [113]. Once properly discretized, the combined regularization $\mathcal{L}_{PLN}$ yields a 37.2% NMSE improvement at 1–10% sampling rates over diffusion-only baselines [56].
@@ VIII-D3 Physics-Informed Generative Architectures
VIII-D4 Physical Hallucinations: Definition and Detection
Figure 7: Open research challenges and future directions for learning-based radio map construction, organized along three axes: complex propagation scenarios, advanced inversion and optimization, and foundation models with unified frameworks.
VIII-E Summary and Practitioner Guidance
@@ IX Open Issues and Future Directions
Learning-based RM construction has achieved significant milestones [152, 120]. However, the transition toward ubiquitous 6G environmental awareness remains constrained by several open challenges. These span two axes: modeling of uncharted propagation scenarios and algorithmic frontiers in inverse problems and physics-data co-design. This section identifies the most pressing issues, drawing upon the architectural insights, physical embedding strategies, and dataset limitations discussed throughout this tutorial. Fig. 7 organizes these challenges along three complementary axes.
@@ X Concluding Remarks
This tutorial has systematically surveyed the data foundations, neural architectures, and physics-informed strategies for learning-based RM construction. We have established a unified taxonomy organized by the forward-inverse problem dichotomy and a three-level physics integration framework spanning data, loss, and architecture. The overarching lesson is clear: the progression from purely data-driven models toward physics-data dual-driven frameworks is a necessity dictated by the information-theoretic structure of the problem. When observations are sparse and propagation is complex, data alone cannot uniquely determine the RM. The model must draw upon physical priors to resolve ambiguity. The three levels of physics integration surveyed in Section VIII provide progressively deeper embedding of electromagnetic knowledge. Input-level feature engineering offers the lowest entry barrier. Loss-level PDE regularization constrains solutions to feasible electromagnetic manifolds. Architecture-level structural isomorphism achieves the deepest alignment with wave propagation dynamics. Each level trades computational cost for physical fidelity. These advances collectively strengthen downstream tasks such as resource allocation, interference management, and predictive network planning. Several converging trends point toward the future of the field. Differentiable ray tracing enables joint simulation-optimization. Foundation model pre-training supports cross-environment generalization. Amortized inference can bring real-time rendering to optics-inspired methods. Formal hallucination detection provides deployment assurance. Together, these trends suggest a future where RM construction transitions from an offline, per-scenario engineering task to an always-on, physically rigorous component of the wireless network infrastructure. Achieving this vision requires sustained collaboration between the wireless communications, computational electromagnetics, and machine learning communities. The most powerful RM models will be those that respect the physics they seek to represent.
This tutorial has systematically surveyed the data foundations, neural architectures, and physics-informed strategies for learning-based radio map construction. We have established a unified taxonomy organized by the forward-inverse problem dichotomy and introduced a three-level physics integration framework spanning data, loss, and architecture. These advances collectively strengthen downstream tasks in wireless communication networks, including resource allocation, interference management, and predictive network planning, by enabling accurate and efficient radio environment digitization. The overarching lesson is that the progression from purely data-driven models toward physics-data dual-driven frameworks is a necessity dictated by the information-theoretic structure of the problem. When observations are sparse and propagation is complex, data alone cannot uniquely determine the radio map, and the model must draw upon physical priors to resolve ambiguity. Input-level feature engineering offers the lowest entry barrier, loss-level PDE regularization constrains solutions to feasible electromagnetic manifolds, and architecture-level structural isomorphism achieves the deepest alignment with wave propagation dynamics, where each level trades computational cost for physical fidelity. Looking ahead, several converging trends point toward a future where radio map construction transitions from an offline, per-scenario engineering task to an always-on, physically rigorous component of wireless network infrastructure: differentiable ray tracing enables joint simulation-optimization, foundation model pre-training supports cross-environment generalization, amortized inference brings real-time rendering to optics-inspired methods, and formal hallucination detection provides deployment assurance. Achieving this vision requires sustained collaboration between the wireless communications, computational electromagnetics, and machine learning communities, and the most powerful radio map models will ultimately be those that respect the physics they seek to represent.